If the speed of a wave on a violin string is 400 m/s, what will be the speed of the wave if the tension in the string is halved?
Answer #1

here,

Speed of wave is directly proportional to square root of Tension

So

v1/v2 = sqrt(T1/T2)

According to question T2 = T1/2 --> T1 = 2T2

400/v2 = sqrt(2 * T2/T2)

v2 = 400/sqrt(2)

v2 = 283.7 m/s

the speed of the wave if the tension in the string is halved is 283.7 m/s